Please help.
Sky divers. Bob and Betty simultaneously jump from two airplanes at different altitudes. Bob’s altitude t seconds after leaving the plane is -16t^2 + 6600 feet. Betty’s altitude t seconds after leaving the plane is -16t^2 + 7400 feet. Write a polynomial that represents the difference between their altitudes t seconds after leaving the planes. What is the difference between their altitudes 3 seconds after leaving the planes?

Answer by stanbon(75887)

Bob’s altitude t seconds after leaving the plane is -16t^2 + 6600 feet. Betty’s altitude t seconds after leaving the plane is -16t^2 + 7400 feet. Write a polynomial that represents the difference between their altitudes t seconds after leaving the planes. What is the difference between their altitudes 3 seconds after leaving the planes?
------------
Bob's altitude after 3 seconds:
a(t) = -16t^2+6600 =
----------
Betty's altitude after 3 seconds:
b(t) = -16t^2+ 7400
--------
Difference polynomial
h(t) = b(t)-a(t) = 800 ft
--------------
After 3 seconds the difference is 800 ft.
It continues to be 800 ft. till Bob hits the ground or releases his parachute.
--------------
Cheers,
Stan H.